A gas sample occupies a volume of 12.0 L at 50 ̊C and 700 torr. How

many moles of gas are contained in the sample?


1
Expert's answer
2021-09-28T01:04:20-0400

"PV= nRT"

12×700 =n × 8.314×323

n= 8400/2685.422

n =3.13moles